Recent Log Entries

The following cavity histories have been updated: 2L21-6

Log Number: 3998296Logbook(s): Tags: Body: 

The following cavity histories have been updated:
2L21-6
Username: evonda
Date/Time: 06-01-22_17:17:36
RF Status: Bad
Tuner Status: Good
Comment: Heater cable that runs from the card cage to the interconnect board replaced resolving initial issue. Now the cavity is showing KMAI faults will T/S tomorrow. Techs: E. Valentine & S. Holben
Special Instructions:

2L21-6 KMAI Fault dropping HV

Log Number: 3998297Logbook(s): Entry Makers: kernsBody: 

Jumpers put back on cav.6 and 7. Cavity was bypassed for KMAI fault but jumper not pulled and zone won't go to HV. E. Valentine and S. Holben. Work to continue !

Electric heat bottomed out in 2L24/25

Log Number: 3998295Logbook(s): Entry Makers: michaelmBody: 

We may want to raise the electric heat to over 200W in 2L24/25 if the zones can maintain liquid level there. At present they have bottomed out, and there is potentially still some gradient available for use there if the load is compensated.

The following cavity histories have been updated: 2L21-7

Log Number: 3998293Logbook(s): Tags: Body: 

The following cavity histories have been updated:
2L21-7
Username: evonda
Date/Time: 06-01-22_17:15:31
RF Status: Good
Tuner Status: Good
Comment: Jumper for 50 ohm resistor installed and cavity back up and running. Techs: E. Valentine & S. Holben
Special Instructions:

iocsoftagileturbo restored

Log Number: 3998280Logbook(s): Tags: Body: restorelib 2.5 Target restore date was 2022-06-01 16:59:37 The output from the myRestore execution was as follows: iocsoftagileturbo systemEnum@2022-06-01 16:59 2(2) restored Execute: /usr/csite/certified/bin/ced2Epics -ced ced -A -s -i iocsoftagileturbo (success) Procedure return status: Success

2L11 Cavity one and two

Log Number: 3998253Logbook(s): Entry Makers: kernsBody: 

With cav. 2 running for the last three hours, you can see the detune angle start to jump around. Cavity 1 was running at 0500 this morning ~6.8 Mv/m and vacuum took a hit at 0600 when gradient was increased to 7.8 Mv/m. TDETA was stable on cavity 1 when both were on at the same time .

iocsoftagileturbo restored

Log Number: 3998218Logbook(s): Tags: Body: restorelib 2.5 Target restore date was 2022-06-01 16:16:11 The output from the myRestore execution was as follows: iocsoftagileturbo systemEnum@2022-06-01 16:16 2(2) restored Execute: /usr/csite/certified/bin/ced2Epics -ced ced -A -s -i iocsoftagileturbo (success) Procedure return status: Success

Operator SIGN IN

Log Number: 3998192Logbook(s): Body:  ACCELERATOR OPERATOR SIGN IN FORM OPERATOR:    Schaumloffel
DATE:             06/01/2022
SHIFT:               Swing

The operator has read the following:

1L03-4 No CRRP/no gradient

Log Number: 3998182Logbook(s): Entry Makers: kernsBody: 

Circulator really bad and needs rebuilt load installed. After RFCM replaced, power checked, and all input cables checked, still could not get any gradient or CRRP. Raised CRFP up to 3kw and only got 330w CRRP before tripped on KRRP at 480w. Also found a broken pin on the 3db pad probe attenuator which I don't remember replacing but worked on 17 May. Scrolling back to 18 May 22, you can see CRFP at 5.58kw and KRRP spike at 515w for the last time.

Pages